Calculate the energy of a single photon of light with a frequency of 3.00×1016 s-1.
1
Expert's answer
2020-09-23T04:43:36-0400

The energy of a single photon can be calculated using Planck equation:

E = hf

where E - photon energy, h - Planck constant (h = 6.626 × 10-34 J s ), f - photon frequancy

From here:

E = 6.626 × 10-34 J s × 3.00 × 1016 s-1 = 19.878 × 10-18 J = 1.99 × 10-17 J


Answer: 1.99 × 10-17 J